# The impact of pressure oscillations on bubble rising in shear-thinning fluids

### v1

@@ Abstract.
We study the rising dynamics of a bubble driven into periodic volumetric oscillations by an external pressure driving within a highly viscous shear-thinning fluid. We perform axisymmetric direct numerical simulations employing the Carreau-Yasuda model to describe the rheological behavior of the fluid. We carry out a parametric study of the bubble rising dynamics, changing the amplitude and the frequency of the external pressure driving, and the bubble radius. Due to the external pressure oscillations, the bubble undergoes volume changes that strain the liquid at much larger rates than those due to natural rising, causing the surrounding fluid viscosity to thin. The numerical results show that the rising dynamics become highly nonlinear and unsteady due to the interplay of the shear-thinning rheology and the external driving. As a result, the period-averaged rising velocity of the bubble can increase by orders of magnitude compared to its natural rising velocity. These nonlinear effects become progressively more important as the amplitude and the frequency of the pressure driving or the bubble radius are increased. Qualitatively, the simulation model agrees with previous experimental findings in terms of average rising velocity. However, the experiments exhibit terminal velocities that are smaller than those predicted numerically, along with differences in bubble shape during the ascent. These discrepancies may be attributed to modeling the fluid rheology as a generalized Newtonian fluid rather than as a viscoelastic one.
@@ 1.  Introduction
Given its relevance in several industrial applications, it is no surprise that the rising of bubbles in complex fluids has a long history of intense research [20, 34]. Over the past two decades, a combination of high-speed experiments and detailed simulations has revealed how the non-Newtonian fluid rheology impacts the rising of bubbles. Pillapakkam et. al demonstrated that rising bubbles in viscoelastic polymer solutions develop cusp-shaped rear profiles and exhibit a six-fold velocity jump at a critical size. The shape of the bubble and of its wake structure changes fundamentally at the onset of the jump [4].
They also proposed a universal correlation of non-dimensional numbers for the dimensionless critical bubble volume at the jump discontinuity [5]. Subsequently, a computational and theoretical study by Fraggedakis et. al investigated the critical bubble volume in viscoelastic fluids. They found that the velocity discontinuity is due to a hysteresis loop, arising from the formation of a tip at the rear pole. They also found the characteristic negative wake effect, already observed by Hassager[34]. Their predictions achieved quantitative agreement with the experimental results of Pilz and Brenn, as well as qualitative agreement with larger bubbles involving inertia [9].
The rising of millimiter to micrometer-sized bubbles in complex fluids can be extremely slow due to their large viscosity or even impossible if the bubble buoyancy does not overcome the yield stress [8, 36, 35]. One promising approach to remove these bubbles involves enhancing their rising velocity through periodic pressure changes or vibrations. The phenomenon of degassing by acoustic cavitation has been studied for several years [2]. This approach harnesses violent bubble cavitation under strong acoustic fields. Instead, in the case of complex fluids, stimuli at frequencies and amplitudes much lower than that of the cavitation threshold can locally alter viscosity around the bubble or cause the material to yield. This strategy directly leverages the mechanical stress generated by oscillating bubbles to facilitate bubble release. Additionally, ultrasound-induced secondary Bjerknes forces have been suggested as a way to encourage bubble merging, thereby increasing their size and ascent speed [38]. In contrast to chemical antifoaming agents, the application of these mild mechanical stimuli leaves no residual additives, preserving purity and rheology while eliminating concerns about fouling or compatibility.
@@
To assess the importance of shear thinning, we investigated the drag force experienced by a slowly pulsating bubble [11]. In this model, we assumed low-Reynolds numbers, we fixed the kinematics of the bubble and we employed the Carreau–Yasuda model to describe the fluid rheology. We found that the radial oscillations significantly decreased the bubble drag. However, the effect was larger than that observed in the experiments. Zhang et. al solved the unsteady Stokes problem for a spherical bubble translating through a shear-thinning fluid at a constant velocity plus a high-frequency sinusoidal oscillation. They find a transition from quadratic to power-law scaling between drag reduction and oscillation amplitude [13]. Despite their valuable insights, these theories retain restrictive assumptions: fixed spherical shape, negligible inertia, or a clear separation of pulsation and translation timescales. In this work, we study bubble rising in a shear-thinning fluid under an externally-applied oscillatory pressure by solving the fully coupled Navier–Stokes equations for a Carreau–Yasuda fluid. We track the deformable gas–liquid interface of the bubble using an arbitrary Lagrangian–Eulerian (ALE)[19] formulation and account for fluid inertia. This enables the model to capture transient inertial effects and arbitrary bubble deformation.
@@ 2.  Governing equations
To solve the governing equations (1)-(10), we use the finite element method that we developed and validated previously. To fulfill the inf-sup condition, we use quadratic interpolation functions for the velocity field and linear interpolation functions for the pressure. We discretize the axisymmetric computational domain, $\Omega$ in Figure 1, into triangular elements with a more refined mesh close to the bubble surface (See Figure 3). To handle the mesh deformation due to the movements of the bubble surface, we employ the Arbitrary Lagrangian Eulerian (ALE) method[19]. The velocity of the mesh, $\mathbf{v}_{m}$, at any point in the computational domain is obtained by solving a Laplace equation $\nabla^{2}\mathbf{v}_{m}=\mathbf{0}$, with Dirichlet boundary conditions $\mathbf{v}_{m}=\mathbf{v}\cdot\mathbf{n}\mathbf{n}$ on the bubble boundary, $\Gamma_{1}$, and $\mathbf{v}_{m}=dz_{\text{cm}}/dt\,\,\hat{\mathbf{z}}$ on $\Gamma_{2}$. The boundary condition on $\Gamma_{2}$ moves the entire computational boundary upwards whenever the bubble center of mass is displaced. This condition minimizes mesh deformation and removes the need to remesh. We use an implicit time-integration method with an adaptive second-order implicit Euler method. The resulting nonlinear system of equations is solved using the Newton method with a tolerance set to $10^{-8}$ and the direct solver PARDISO.
@@ 2.1.  Characteristic dimensionless numbers
Using the parameters shown in Table I, and the range of bubble radii $R\in[0.5,2\rm{mm}]$, the Bond number spans the range $Bo\in[0.0324,0.5190]$. As a consequence, the bubble equilibrium shape is very close to a sphere. This agrees with the experimental observations[3][1].
@@
To verify this point, we performed numerical simulations of bubble rising in the absence of a pressure driving by considering a no-slip boundary condition on $\Gamma_{2}$. We computed the steady-state natural rising speed of the bubbles, $v_{0}$, which are reported in Table II. In the case of $R=1\rm{mm}$, the rising velocities of the bubbles, $v_{0}$ are close to those predicted by the Stokes law for a bubble, $v_{\text{Stokes}}=\rho gR^{2}/3\eta_{0}$ [3]. As the bubble becomes bigger the natural rising velocity deviates by a factor of up 2.5 in the case of the largest bubble radius $R=2\rm{mm}$ due to increasingly large shear-thinning of the fluid surrounding the bubble. The values of the natural rising velocity of the bubbles, $v_{0}$, reported in Table II will be used as a comparison for the rising speed of bubbles driven by pressure oscillations.
@@
$5\text{mm} 0.0$104 0.0091
$1\text{mm}$ 0.0534 0.0363
$2\text{mm}$ 0.4135 0.1453
@@
In summary, in the absence of pressure oscillations, we expect the bubbles considered in this work to rise at a speed close to that predicted by the Stokes law and to remain close to its initial spherical shape. However, when the pressure oscillations are applied the bubble experiences fast volumetric oscillations which considerable change the viscosity and its rising speed. Since we consider pressure oscillating at frequencies in the range of 1 to 300 Hz, the Carreau number based on the frequency $Cu_{\omega}$, given
@@ 4.  Results
It is useful to compare our numerical results to those of Iwata et al[1] at $f=300\rm{Hz}$. In Figure 6(a-b), we report the evolution of the bubble vertical and horizontal diameter measured by Iwata et al. and in our simulations. In our simulations, the vertical diameter is computed as the difference between the largest and smallest z-coordinate of the bubble surface, and the horizontal diameter as twice the largest r-coordinate of the bubble surface (see Figure 6(e-h)). We considered a 1mm diameter bubble, which is slightly smaller than that used in the experiments. Since in the experiments they do not control the pressure amplitude, we had to guess the value of $k$ that adjusts to the experiments. By using a value of $k=0.75$, we find that the excursions of the bubble vertical and horizontal diameters in the experiments and in the simulations are similar. In our simulations, the bubble remains almost spherical throughout the compression-expansion cycle. This is shown by the overlapping green and blue curves in Figure 6(b) and it is also apparent in the snapshots, Figure 6(e-h). During most of the compression and expansion of the bubble, our numerical results are in qualitative agreement with the experiments, Figure 6(20,4,15)). However, close to the minimum radius of the bubble Figure 6(24)), the bubble develops a weak cusp at the rear, which is the signature of viscoelastic effects that are ignored in our model. The onset of the cusp is also apparent in Figure 6(a), whereby the vertical and horizontal diameters in the experiments do not overlap anymore.
Figure 6. Comparison between our numerical results and the experiments of Iwata et al.[1] Panel (a) shows the horizontal and vertical measurements in the experiments. Panel (b) shows the same quantities obtained in the numerical simulations. Panels (e-h) show a sequence of bubble snapshots at different time instants for the case $(f,k,R_{0})=(300~\text{Hz},0.75,0.5~\text{mm})$. The right panels (4,15,20,24) show reference snapshots taken directly from Iwata et al. [1], with time indices corresponding to specific points in the graph (Reproduced with permission from J. non-Newton. Fluid Mech. 151, (1-3) (2008) Copyright Elsevier 2008).
@@
Figure 7. Kinematics of the bubble for the case $R_{0}=0.5\ \text{mm}$. The dimensionless time is scaled with $T=1/f$ and $(f_{1},f_{4},f_{5})=(1,150,300)\ \text{Hz}$. Note that the vertical axes have different ranges in the three panels.
@@
Figure 8. Fourier analysis of the velocity and volume of the bubble normalized by their mean value. We consider the cases of two amplitudes $k=0.05$ and $k=0.2$, with $R_{0}=0.5\text{mm}$ and $f=75\text{Hz}$.
@@
Figure 10. Analysis of the dimensionless numbers showing the variation of the Reynolds number (calculated from eq. 16) and the Womersley number (calculated from equation eq. 17)) for different amplitudes and frequencies. Each panel corresponds to a different bubble radius, and each marker indicates a specific driving amplitude.
@@ 4.1.  Comparison with Iwata Experiments
Previous experimental studies [1] have focused on the rising of millimetric bubbles subjected to a periodically oscillating pressure. In this study, we compare our simulation results with these experimental findings, focusing on bubble shape and terminal rise velocity. The reference study conducted experiments within a system referred to as POD (Pressure Oscillation Defoaming), consisting of a quartz cell filled with the sample liquid and sealed with a rubber diaphragm cap. A vibrating piston, driven by a sinusoidal generator, induced periodic pressure oscillations inside the cell, causing the bubble to undergo repeated compression and expansion. These oscillations generated local radial flows, which, for shear-thinning fluids, reduced the apparent viscosity around the bubble and enhanced its rise velocity. The rising dynamics of the bubbles were analyzed as a function of a non-dimensional radial acceleration of their surface, $G_{b}$, defined as $G_{b}=(2\,\pi f)^{2}(D_{\text{max}}-D_{\text{min}})/4g$ where $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum diameter attained by the bubble during one driving period, respectively. The dimensions of their bubbles are comparable to those considered in the present study. Figure 11 shows their experimental data overlaid with the numerical results, where the x-axis represents the non-dimensional acceleration and the y-axis represents the velocity normalized by the no-oscillation case.
@@
The study by Iwata et. al [1] did not report the amplitude of the oscillations imposed on the system, and our investigation does not identify a value capable of fitting their data. In addition, we also observe discrepancies with the experiments in the bubble shape during a driving period and in the trend of the average rising velocity as a function of the acceleration. The primary reason may be that a shear-thinning model alone cannot capture the full rheology, since the fluid used in their experiments may exhibit elastic effects. Elastic stresses combined with shear thinning can explain the formation of the cusp observed in Figure 6 [8, 27]. Further simulations will be conducted considering constitutive models that account for fluid elasticity.
Figure 11. Comparison of simulated and experimental results. The values in millimeters refer to the initial radius of the various simulations. The x-axis is a parameter introduced by Iwata et al. [1], where $D_{\text{Max}}$ and $D_{\text{Min}}$ are the maximum and minimum diameters that the bubble reaches during the rise.

### v2

@@ Abstract.
We study the rising dynamics of a bubble driven into periodic volumetric oscillations by an external pressure driving within a highly viscous shear-thinning fluid. We perform axisymmetric direct numerical simulations employing the Carreau-Yasuda model to describe the rheological behavior of the fluid and the finite element method to discretize the equations. We carry out a parametric study of the bubble rising dynamics, changing the amplitude and the frequency of the external pressure driving, and the bubble radius. Due to the external pressure oscillations, the bubble undergoes volume changes that strain the liquid at much larger rates than those due to natural rising, causing the surrounding fluid viscosity to thin. The numerical results show that the rising dynamics become highly nonlinear and unsteady due to the interplay of the shear-thinning rheology and the external driving. As a result, the period-averaged rising velocity of the bubble can increase by orders of magnitude compared to its natural rising velocity. These nonlinear effects become progressively more important as the amplitude and the frequency of the pressure driving or the bubble radius are increased. Qualitatively, the simulation model agrees with previous experimental findings in terms of average rising velocity. However, the experiments exhibit terminal velocities that are smaller than those predicted numerically, along with differences in bubble shape during the ascent. These discrepancies may be attributed to modeling the fluid rheology as a generalized Newtonian fluid rather than as a viscoelastic one.
@@ 1. Introduction
Given its relevance in several industrial applications, it is no surprise that the rising of bubbles in complex fluids has a long history of intense research [20, 34]. Over the past two decades, a combination of high-speed experiments and detailed simulations has revealed how the non-Newtonian fluid rheology impacts the rising of bubbles. Pillapakkam et. al demonstrated that rising bubbles in viscoelastic polymer solutions develop cusp-shaped rear profiles and exhibit a six-fold velocity jump at a critical size. The shape of the bubble and of its wake structure changes fundamentally at the onset of the jump [4]. They also proposed a universal correlation of non-dimensional numbers for the dimensionless critical bubble volume at the jump discontinuity [5]. Subsequently, a computational and theoretical study by Fraggedakis et. al investigated the critical bubble volume in viscoelastic fluids. They found that the velocity discontinuity is due to a hysteresis loop, arising from the formation of a tip at the rear pole. They also found the characteristic negative wake effect, already observed by Hassager[34]. Their predictions achieved quantitative agreement with the experimental results of Pilz and Brenn, as well as qualitative agreement with larger bubbles involving inertia [9].
The rising of millimiter to micrometer-sized bubbles in complex fluids can be extremely slow due to their large viscosity or even impossible if the bubble buoyancy does not overcome the yield stress [8, 36, 35, 40]. One promising approach to remove these bubbles involves enhancing their rising velocity through periodic pressure changes or vibrations. The phenomenon of degassing by acoustic cavitation has been studied for several years [2]. This approach harnesses violent bubble cavitation under strong acoustic fields. Instead, in the case of complex fluids, stimuli at frequencies and amplitudes much lower than that of the cavitation threshold can locally alter viscosity around the bubble or cause the material to yield. This strategy directly leverages the mechanical stress generated by oscillating bubbles to facilitate bubble release. Additionally, ultrasound-induced secondary Bjerknes forces have been suggested as a way to encourage bubble merging, thereby increasing their size and ascent speed [38]. In contrast to chemical antifoaming agents, the application of these mild mechanical stimuli leaves no residual additives, preserving purity and rheology while eliminating concerns about fouling or compatibility.
@@
To assess the importance of shear thinning, we investigated the drag force experienced by a slowly pulsating bubble [11]. In this model, we assumed low-Reynolds numbers, we fixed the kinematics of the bubble and we employed the Carreau–Yasuda model to describe the fluid rheology. We found that the radial oscillations significantly decreased the bubble drag. However, the effect was larger than that observed in the experiments.
Zhang et. al solved the unsteady Stokes problem for a spherical bubble translating through a shear-thinning fluid at a constant velocity plus a high-frequency sinusoidal oscillation. They find a transition from quadratic to power-law scaling between drag reduction and oscillation amplitude [13].
Despite their valuable insights, these theories retain restrictive assumptions: fixed spherical shape, negligible inertia, or a clear separation of pulsation and translation timescales. Such assumptions are removed in the present work, which addresses a key research gap left by previous studies. Here, we study bubble rising in a shear-thinning fluid under an externally-applied oscillatory pressure by solving the fully coupled Navier–Stokes equations for a Carreau–Yasuda fluid. We track the deformable gas–liquid interface of the bubble using an arbitrary Lagrangian–Eulerian (ALE)[19] formulation and account for fluid inertia. This enables the model to capture transient inertial effects and arbitrary bubble deformation.
@@ 2. Governing equations
Table 1. Fluid parameters used in the simulations
Physical property Value and unit
$\rho$ $1003.5~\mathrm{kg/m^{3}}$
$\sigma$ $0.072~\mathrm{N/m}$
$\eta_{0}$ $90~\mathrm{Pa\cdot s}$
$\eta_{\infty}$ $10^{-3}~\mathrm{Pa\cdot s}$
$a$ 2
$n$ 0.3
$\lambda$ 57 s
To solve the governing equations (1)-(10), we use the finite element method that we developed and validated previously. The finite element method has been widely employed to study the dynamics of bubbles rising in complex fluids [9]. To fulfill the inf-sup condition, we use quadratic interpolation functions for the velocity field and linear interpolation functions for the pressure. We discretize the axisymmetric computational domain, $\Omega$ in Figure 1, into triangular elements with a more refined mesh close to the bubble surface (See Figure 3). To handle the mesh deformation due to the movements of the bubble surface, we employ the Arbitrary Lagrangian Eulerian (ALE) method[19]. The velocity of the mesh, $\mathbf{v}_{m}$, at any point in the computational domain is obtained by solving a Laplace equation $\nabla^{2}\mathbf{v}_{m}=\mathbf{0}$, with Dirichlet boundary conditions $\mathbf{v}_{m}=\mathbf{v}\cdot\mathbf{n}\mathbf{n}$ on the bubble boundary, $\Gamma_{1}$, and $\mathbf{v}_{m}=dz_{\text{cm}}/dt\,\,\hat{\mathbf{z}}$ on $\Gamma_{2}$. The boundary condition on $\Gamma_{2}$ moves the entire computational boundary upwards whenever the bubble center of mass is displaced. This condition minimizes mesh deformation and removes the need to remesh. We use an implicit time-integration method with an adaptive second-order implicit Euler method. The resulting nonlinear system of equations is solved using the Newton method with a tolerance set to $10^{-8}$ and the direct solver PARDISO.
@@ 2.1. Characteristic dimensionless numbers
Using the parameters shown in Table I, and the range of bubble radii $R\in[0.5,2]\,\mathrm{mm}$, the Bond number spans the range $Bo\in[0.0324,0.5190]$. As a consequence, the bubble equilibrium shape is very close to a sphere. This agrees with the experimental observations [3, 1].
@@
To verify this point, we performed numerical simulations of bubble rising in the absence of a pressure driving by considering a no-slip boundary condition on $\Gamma_{2}$. We computed the steady-state natural rising speed of the bubbles, $v_{0}$, which are reported in Table II. In the case of $R=1\,\mathrm{mm}$, the rising velocities of the bubbles, $v_{0}$, are close to those predicted by the Stokes law for a bubble, $v_{\text{Stokes}}=\rho gR^{2}/3\eta_{0}$ [3]. As the bubble becomes bigger, the natural rising velocity deviates by a factor of up 2.5 in the case of the largest bubble radius $R=2\,\mathrm{mm}$ due to increasingly large shear-thinning of the fluid surrounding the bubble. The values of the natural rising velocity of the bubbles, $v_{0}$, reported in Table II will be used as a comparison for the rising speed of bubbles driven by pressure oscillations.
@@
$5\,\mathrm{mm} 0.0$104 0.0091
$1\,\mathrm{mm}$ 0.0534 0.0363
$2\,\mathrm{mm}$ 0.4135 0.1453
@@
In summary, in the absence of pressure oscillations, we expect the bubbles considered in this work to rise at a speed close to that predicted by the Stokes law and to remain close to its initial spherical shape. However, when the pressure oscillations are applied, the bubble experiences fast volumetric oscillations which considerably change the viscosity and its rising speed. Since we consider pressure oscillating at frequencies in the range of 1 to 300 Hz, the Carreau number based on the frequency $Cu_{\omega}$, given by
@@ 4. Results
It is useful to compare our numerical results to those of Iwata et al[1] at $f=300\rm{Hz}$. In Figure 6(e-f), we report the evolution of the bubble vertical and horizontal diameter measured by Iwata et al. and in our simulations. In our simulations, the vertical diameter is computed as the difference between the largest and smallest z-coordinate of the bubble surface, and the horizontal diameter as twice the largest r-coordinate of the bubble surface (see Figure 6(a-d)). We considered a 1mm diameter bubble, which is slightly smaller than that used in the experiments. Since in the experiments they do not control the pressure amplitude, we had to guess the value of $k$ that adjusts to the experiments. By using a value of $k=0.75$, we find that the excursions of the bubble vertical and horizontal diameters in the experiments and in the simulations are similar. In our simulations, the bubble remains almost spherical throughout the compression-expansion cycle. This is shown by the overlapping green and blue curves in Figure 6(f) and it is also apparent in the snapshots Figure 6(a-d). During most of the compression and expansion of the bubble, our numerical results are in qualitative agreement with the experiments, Figure 6(e-f)). However, close to the minimum radius of the bubble Figure 6(e)), the bubble develops a weak cusp at the rear, which is the signature of viscoelastic effects that are ignored in our model. The onset of the cusp is also apparent in Figure 6(i), whereby the vertical and horizontal diameters in the experiments do not overlap anymore.
Figure 6. Comparison between our numerical results and the experiments of Iwata et al.[1] Panel (a) shows the horizontal and vertical measurements in the experiments. Panel (b) shows the same quantities obtained in the numerical simulations. Panels (e-h) show a sequence of bubble snapshots at different time instants for the case $(f,k,R_{0})=(300~\text{Hz},0.75,0.5~\text{mm})$. The right panels (g-l) show reference snapshots taken directly from Iwata et al. [1], with time indices corresponding to specific points in the graph (Reproduced with permission from J. non-Newton. Fluid Mech. 151, (1-3) (2008) Copyright Elsevier 2008).
@@
Figure 7. Kinematics of a bubble with initial radius $R_{0}=0.5~\text{mm}$ under different forcing amplitudes $k=0.05,0.2,0.75$. Dimensionless time tftf is used, scaled with forcing frequencies $(f_{1},f_{4},f_{5})=(1,150,300)~\text{Hz}$. Panels (a-c) show the dimensionless vertical position of the bubble $Z_{b}(t)/R_{0}$ for increasing forcing frequencies: (a) $f_{1}=1~\text{Hz}$, (b) $f_{4}=150~\text{Hz}$, and (c) $f_{5}=300~\text{Hz}$. Panels (d-f) report the corresponding dimensionless bubble velocity $v_{b}(t)/v_{0}$ for the same frequencies. Note that the vertical axes have different ranges in the three panels.
@@
Figure 8. Fourier analysis of (a) the velocity and (b) the volume of the bubble normalized by their mean value. We consider the cases of two amplitudes $k=0.05$ and $k=0.2$, with $R_{0}=0.5\text{mm}$ and $f=75\text{Hz}$.
@@
Figure 10. Analysis of the dimensionless numbers: the Reynolds number $\mathrm{Re}$ (computed from (16)) and the Womersley number $\mathrm{Wo}$ (computed from (17)), shown as functions of forcing amplitude and frequency. Panels (a-c) correspond to fixed initial bubble radii $R_{0}=0.5\ \mathrm{mm}$ (a), $R_{0}=1\ \mathrm{mm}$ (b) and $R_{0}=2\ \mathrm{mm}$ (c), respectively. Each marker indicates a specific driving amplitude; different marker styles denote different driving frequencies (see legend).
@@ 4.1. Comparison with Iwata Experiments
Previous experimental studies [1] have focused on the rising of millimetric bubbles subjected to a periodically oscillating pressure. In this study, we compare our simulation results with these experimental findings, focusing on bubble shape and terminal rise velocity. The reference study conducted experiments within a system referred to as POD (Pressure Oscillation Defoaming), consisting of a quartz cell filled with the sample liquid and sealed with a rubber diaphragm cap. A vibrating piston, driven by a sinusoidal generator, induced periodic pressure oscillations inside the cell, causing the bubble to undergo repeated compression and expansion. These oscillations generated local radial flows, which, for shear-thinning fluids, reduced the apparent viscosity around the bubble and enhanced its rise velocity. The rising dynamics of the bubbles were analyzed as a function of a non-dimensional radial acceleration of their surface, $G_{b}$, defined as
$G_{b}=\frac{(2\pi f)^{2}\bigl(D_{\text{max}}-D_{\text{min}}\bigr)}{4g},$
where $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum diameters attained by the bubble during one driving period, respectively. The dimensions of their bubbles are comparable to those considered in the present study. Figure 11 shows their experimental data overlaid with the numerical results, where the x-axis represents the non-dimensional acceleration and the y-axis represents the velocity normalized by the no-oscillation case.
@@
The study by Iwata et al. [1] did not report the amplitude of the oscillations imposed on the system, and our investigation does not identify a value capable of fitting their data. In addition, we also observe discrepancies with the experiments in the bubble shape during a driving period and in the trend of the average rising velocity as a function of the acceleration. The primary reason may be that a shear-thinning model alone cannot capture the full rheology, since the fluid used in their experiments may exhibit elastic effects. Elastic stresses combined with shear thinning can explain the formation of the cusp observed in Figure 6 [8, 27]. Further simulations will be conducted considering constitutive models that account for fluid elasticity.
Figure 11. Comparison of simulated and experimental results. The values in millimeters in the legend refer to the initial radius of the bubble in the simulations. The bottom axis was introduced by Iwata et al. [1] and represents a characteristic radial acceleration of the bubble surface. Here $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum diameters that the bubble reaches during the rise. The experimental data points are extracted from Figure 8 of Iwata et al. [1].